OpenStudy (anonymous):

what is the product of the radical expression (7-sqrt2)(8+sqrt2) a. 13+15 sqrt 2 b. 54+56 sqrt 2 c. 54- sqrt 2 d was wrong

OpenStudy (kira_yamato):

(7-sqrt2)(8+sqrt2) = 56 + 7 sqrt(2) - 8 sqrt(2) - 2 = 54 - sqrt(2)

OpenStudy (anonymous):

so its b

OpenStudy (anonymous):

C my bad

OpenStudy (kira_yamato):

Yep

OpenStudy (anonymous):

also do you know how to do this sqrt 2+4 - sqrt x = 2

OpenStudy (anonymous):

@Kira_Yamato

OpenStudy (kira_yamato):

sqrt(2) + 4 - sqrt x = 2 sqrt(2) - sqrt(x) = -2 sqrt(x) = 2 + sqrt(2) x = 6 + 4sqrt(2)

OpenStudy (anonymous):

a.x=0 and x=-16 b.x=0 and x=16 c.x=0 d was wrong

OpenStudy (anonymous):

I think c

OpenStudy (kira_yamato):

Huh?

OpenStudy (anonymous):

a.x=0 and x=-16 b.x=0 and x=16 c.x=0 d was wrong